OpenStudy (anonymous):

1/sin^2x + 1/cos^sx = 1/sin^2x-sin^4x verify identity

hero (hero):

\[\large\frac{1}{\sin^2x} + \frac{1}{\cos^2x} = \frac{1}{\sin^2x - \sin^4x}\] Is this what you have to verify?

OpenStudy (anonymous):

yes

OpenStudy (dumbcow):

just combine fractions .... common denominator is sin^2 cos^2 note Pythagorean identity sin^2 + cos^2 = 1